OpenStudy (anonymous):

How would you rewrite *1000^log(x)* as an expression of x which does not involve any logs?? Help pleaseee :( I got the other 2. I don't know how to get this one

OpenStudy (anonymous):

is this \[1000^{\log(x)}\]?

OpenStudy (anonymous):

yes. the log(x) that is squared, has a base of 10.

OpenStudy (anonymous):

let me try again. write is \[10^{3\log(x)}\] \[(10^{\log(x)})^3\] \[x^3\]

OpenStudy (anonymous):

worked ^^ thanks